# TAIL: Text-Audio Incremental Learning

### v1

@@ 4.1 Experimental Setup
AI-CL [38] 6.96M 14.28 41.13 56.19 17.58 45.50 58.66 5.13 15.31 23.35 4.20 14.35 23.64
VPT (shallow) [20] 3.69M 26.85 63.65 78.23 37.45 70.34 82.68 9.53 26.18 37.07 9.67 25.84 38.18
VPT (deep) [20] 3.69M 20.35 51.38 67.52 25.23 54.63 70.34 7.00 21.21 31.29 6.12 18.28 29.57
I-VL [23] 3.68M 28.64 64.19 78.87 36.24 69.26 84.03 10.51 26.10 36.86 10.62 25.55 36.36
LoRA [10] 4.75 M 28.00 62.77 76.97 35.97 66.85 79.73 9.80 26.26 37.97 8.33 24.69 35.89
@@
Method Parameters BBC Sound Effects (Trained on Step 3) Audioset (Current Step)
@@
AI-CL [38] 6.96M 2.50 5.85 8.90 2.25 9.50 15.75 3.90 9.40 12.87 7.00 20.67 29.33
VPT (shallow) [20] 3.69M 2.75 7.55 10.95 3.00 16.00 24.50 6.80 13.27 16.23 15.00 35.67 48.17
VPT (deep) [20] 3.69M 2.65 7.80 10.70 3.25 13.25 23.00 5.57 12.33 15.53 11.50 31.33 43.50
I-VL [23] 3.68M 2.50 8.05 10.95 3.75 17.00 25.00 6.90 13.53 16.43 16.83 39.50 50.33
LoRA  [10] 4.75M 3.00 8.05 11.25 4.50 15.75 23.25 6.63 13.73 17.07 17.17 36.67 48.17
@@
Table 2: Results of Average Retrieval across Four datasets. (Here Ft means Finetune).
Method Audio-to-Text Text-to-Audio
R@1 R@5 R@10 R@1 R@5 R@10
AI-CL [38] 6.45 17.92 25.33 7.76 22.50 31.85
VPT (shallow) [20] 11.48 27.66 35.62 16.28 36.96 48.38
VPT (deep) [20] 8.89 23.18 31.26 11.53 29.37 41.60
I-VL [23] 12.14 27.97 35.78 16.86 37.83 48.93
LoRA [10] 11.86 27.70 35.81 16.49 35.99 46.76
Ft (Sequential) 12.64 27.59 35.01 19.16 39.41 50.14
Ft (Joint) 9.80 25.42 33.83 14.19 34.13 46.74
PTAT 13.10 29.69 38.11 19.30 40.35 51.61
Evaluation Metrics. We evaluate the audio-text cross-modal retrieval performance of all methods in our experiments using $Recall@1$, $Recall@5$, and $Recall@10$. For a given query, $Recall@k$ is 1 if the positive item appears within the top $k$ retrieved items, and 0 otherwise. In the context of incremental learning, we also consider the average performance across all seen datasets. The $AverageRecall@k$ of a model is the mean of $Recall@k$ values across each dataset, and is defined as:
$AverageRecall@k=\frac{1}{m}\sum_{m=1}^{M}{Recall@k}_{m}$ (10)
where ${Recall@k}_{m}$ denotes the $Recall@k$ on the $m$-th dataset after completing training on the $m$-th dataset (incremental step $m$).
Implementation Details. We select HT-SAT [6] as the audio encoder and ROBERTA [72] as the text encoder, given their widespread use in previous work [9, 61] and proven effectiveness in audio-text learning. For the HT-SAT audio encoder, all audio clips are randomly cropped or padded to 10 seconds to meet HT-SAT’s fixed-size input requirement.
@@
(a) Step Results for AudioCaps. (AudioCaps is trained in step 1.)
(b) Step Results for Clotho. (Clotho is trained in step 2.)
Figure 4: Incremental step evaluation on AudioCaps dataset (step 1) and Clotho dataset (step 2).
4.2 Baselines
As we are the first to introduce the task of Text-Audio Incremental Learning and there is no prior related work, we selected the state-of-the-art method in the field of audio incremental learning AI-CL [38] and several effective prompt tuning approaches, VPT [20] and I-VL [22], to build a benchmark. We also selected another popular parameter-efficient transfer learning approach LoRA [10], which learns low-rank matrices to approximate parameter updates and reduce the number of trainable parameters. These methods have been shown to be effective in fine-tuning models and enabling efficient transfer across various tasks. Meanwhile, we also compare our method with the fine-tuning method.
@@ 4.2 Baselines
  • AI-CL [38]. We use the Audio Spectrogram Transformer with Frequency-Time factorized Attention, as proposed in their method, as our audio encoder. In step 1, we fully train both the audio encoder and the text encoder; in subsequent steps, we follow their approach. Specifically, we utilize the proposed adapter, training only the adapter while freezing the rest of the structure. The model’s input and training setup are the same as in our method.
  • VPT [20]. We compared the two types of prompts discussed in the paper: (1) VPT (shallow) prompts and (2) VPT (deep) prompts. For the shallow prompt approach, we add an audio prompt before the input to the audio encoder. In contrast, for the deep prompt approach, we add a learnable prompt at each layer of the audio encoder. During training, we only train the prompts and projection layers while freezing the rest of the model.
  • I-VL [22]. We incorporate a text prompt into our text embeddings, while the audio embeddings remain the same. Similar to VPT, we freeze the audio and text encoder, only training the text prompts and projection layers.
  • LoRA [10]. To apply LoRA, we applies trainable low-rank updates to $W_{q}$ and $W_{v}$ in Multi-Head Self Attention, and we’ve integrated this into the audio encoder, freezing its other parameters and the text encoder.
  • Finetuning approach. We designed two variations: (1) Finetune (Sequential). Follows a sequential finetuning process on each dataset, aligning with incremental learning. (2) Finetune (Joint). Employs a joint training approach, combining all datasets. In the two finetuning methods, all parameters are trained without additional prompts or knowledge distillation, effectively using the basic contrastive audio-text learning approach from [9].
@@
Table 3: Ablation study on the effect of adding audio prompts at different layers of the audio encoder.
@@
Table 4: Ablation study on different feature and similarity distillation loss items in our method.
@@
R@1 R@5 R@10 R@1 R@5 R@10
Base 11.52 27.10 35.15 16.33 36.74 47.53
+$\mathcal{L}_{FD}$ 12.42+0.90 28.78+1.68 37.02+1.87 18.19+1.86 38.24+1.50 50.50+2.97
+$\mathcal{L}_{SD}$ 13.02+1.50 29.33+2.23 38.15+3.00 18.14+1.81 39.24+2.50 51.11+3.58
+$\mathcal{L}_{SD}$,$\mathcal{L}_{FD}$ 13.10+1.58 29.69+2.59 38.11+2.96 19.30+2.97 40.35+3.61 50.61+3.08
Table 5: Ablation study on different dataset training sequence.
Method Audio-to-Text Text-to-Audio
R@1 R@5 R@10 R@1 R@5 R@10
VPT(shallow) 9.77 24.77 32.89 13.70 32.44 43.07
LoRA 10.75 25.46 33.59 14.66 33.40 44.22
Ft.(Sequential) 11.77 27.07 34.70 15.64 35.71 44.54
PTAT 12.55 28.20 35.95 18.53 37.95 49.29
4.3 Main Results
We present our main experimental results after sequential training on four datasets, with the training order being AudioCaps(step 1), Clotho (step 2), BBC Sound Effects (step 3), and Audioset (current step). As shown in Tab. 1, Our method PTAT significantly outperforms AI-CL, VPT (shallow), VPT (deep), and I-VL in the 1-3 steps, specifically AudioCaps (step 1) and Clotho (step 2). In the current step, we maintain a comparable level with these four methods in both the Audio-to-Text and Text-to-Audio subtasks across $R@1$, $R@5$, and $R@10$ metrics. This demonstrates that our method more effectively prevents catastrophic forgetting and exhibits superior model generalization capabilities.
@@ 4.3 Main Results
To provide a more comprehensive comparison of model performance, we use the average recall across these four datasets as the final evaluation metric for each method. As shown in Tab. 2, in the Text-to-Audio and Audio-to-Text retrieval tasks, our method outperforms other methods on $R@1$, $R@5$, and $R@10$ metrics. Among methods with similar parameter sizes, including VPT (shallow), VPT (deep), and I-VL, I-VL achieves the best results. Compared to the I-VL method with a similar parameter scale, our method achieves an average improvement of 7% on average. Due to the domain gaps across datasets, the Finetune (Joint) method, despite using a larger parameter size, performs significantly worse than the Finetune (Sequential) method. Compared to the large-scale Finetune (Sequential) method, our method only requires 2.42% of its parameters, while still achieving 4.46% higher performance.
To more intuitively demonstrate the model’s resistance to forgetting, as shown in Fig. 4, we focus on the model’s memory performance on the first two datasets. Fig. 4(a) illustrates the incremental learning performance on AudioCaps (starting training at step 1), while Fig. 4(b) shows the test accuracy of Clotho (starting training at step 2) at each incremental step. It is evident that our PTAT method experiences the smallest accuracy drop in each incremental step and consistently maintains the highest accuracy, demonstrating strong resistance to forgetting and stable accuracy throughout the training process. Additionally, we observed that VPT (deep) exhibits more severe forgetting compared to VPT (shallow), suggesting that due to the relatively simple structure of the audio encoder, excessive prompts may lead to more pronounced forgetting.
@@
4.4 Ablation Studies
@@
Prompts Adding Layer. To analyze the different effects of adding prompts at various layers within the encoder, we conducted experiments on different transformer layers of the audio encoder, with the results shown in Tab. 3. Since HTSAT only has 4 layers, we conducted experiments on the first, second, and third layers. We have found that adding prompts at Layer 1 significantly enhances the model’s resistance to forgetting, particularly achieving great performance on the AudioCaps and Clotho datasets. It also requires fewer parameters than adding prompts at Layer 2 and Layer 3. In contrast, adding prompts at Layer 3 helps the model learn new datasets, as evidenced by improved results on the last two datasets, BBC Sound Effects and Audioset. However, it reduces the model’s ability to retain information from previously trained datasets.
@@
Similarity and Feature Distillation. We analyze the effectiveness of our proposed Audio-Text Similarity and Feature Distillation module by removing single or both parts from our PTAT model. As shown in Tab. 4, using Similarity and Feature Distillation (Full ($\mathcal{L}$)) achieves the best performance compared to no distillation or using only individual Similarity/Feature Distillation.This demonstrates the effectiveness of our proposed Similarity and Feature Distillation module.Among them, the combination of removing both Similarity Distillation ($\mathcal{L}_{SD}$) and Feature Distillation ($\mathcal{L}_{FD}$) performs the worst. Compared to adding only Feature Distillation, adding only Similarity Distillation results in a more significant improvement.
Different Dataset training sequence. To reflect real-world data dynamics, we tested our model’s performance by varying the sequence of dataset additions (AudioCaps trained in step 1, Clotho trained in step 2, BBC Sound Effects trained in step 3, AudioSet trained in step 4). As shown in Fig. 5, the results confirm the effectiveness of our method, regardless of the order of the data set. This will provide insights into how the model can adapt to more flexible data updates.
5 Conclusion
@@
This work introduces TAIL, a novel text-audio incremental learning task. Its challenges focused on achieving continuous text-audio retrieval tasks across sequentially expanding datasets, which is previously unexplored. To benchmark the TAIL task, we propose a parameter-efficient method named PTAT, which tackles the task by learning cross-modal prompts tuning and incorporating an audio-text similarity and feature distillation module.  PTAT primarily comprises two components: Audio-Text Prompt Generation (ATPG) and Audio-Text Similarity and Feature Distillation (ATSFD). ATPG generates a series of interconnected audio and text prompts, leveraging prompt tuning to optimize parameters efficiently. ATSFD mitigates catastrophic forgetting by constraining the output of the new model based on the output of the previous model. We benchmark our method and previous incremental learning or prompt tuning methods on AudioCaps, Clotho, BBC Sound Effects and Audioset datasets, and our method outperforms previous methods significantly, particularly demonstrating stronger resistance to forgetting on older datasets.

### v2

@@ 4.1. Experimental Setup
AI-CL (Muthuchamy Selvaraj et al., 2023) 6.96M 14.28 41.13 56.19 17.58 45.50 58.66 5.13 15.31 23.35 4.20 14.35 23.64
VPT (Shallow) (Jia et al., 2022) 3.69M 26.85 63.65 78.23 37.45 70.34 82.68 9.53 26.18 37.07 9.67 25.84 38.18
VPT (Deep) (Jia et al., 2022) 3.69M 20.35 51.38 67.52 25.23 54.63 70.34 7.00 21.21 31.29 6.12 18.28 29.57
I-VL (Ju et al., 2022a) 3.68M 28.64 64.19 78.87 36.24 69.26 84.03 10.51 26.10 36.86 10.62 25.55 36.36
LoRA (et al., 2022) 4.75 M 28.00 62.77 76.97 35.97 66.85 79.73 9.80 26.26 37.97 8.33 24.69 35.89
@@
Method Parameters BBC Sound Effects (Introduced at Step 3) AudioSet (Introduced at Current Step)
@@
AI-CL (Muthuchamy Selvaraj et al., 2023) 6.96M 2.50 5.85 8.90 2.25 9.50 15.75 3.90 9.40 12.87 7.00 20.67 29.33
VPT (Shallow) (Jia et al., 2022) 3.69M 2.75 7.55 10.95 3.00 16.00 24.50 6.80 13.27 16.23 15.00 35.67 48.17
VPT (Deep) (Jia et al., 2022) 3.69M 2.65 7.80 10.70 3.25 13.25 23.00 5.57 12.33 15.53 11.50 31.33 43.50
I-VL (Ju et al., 2022a) 3.68M 2.50 8.05 10.95 3.75 17.00 25.00 6.90 13.53 16.43 16.83 39.50 50.33
LoRA  (et al., 2022) 4.75M 3.00 8.05 11.25 4.50 15.75 23.25 6.63 13.73 17.07 17.17 36.67 48.17
@@
[…]
(10) $\text{AFS}_{\mathcal{D}}=\frac{\text{Recall@10}^{M}_{\mathcal{D}}}{\text{Recall@10}^{m_{0}}_{\mathcal{D}}}.$
Implementation Details. We select HT-SAT (Chen et al., 2022) as the audio encoder and ROBERTA (Zhuang et al., 2021) as the text encoder, given their widespread use in previous work (Elizalde et al., 2023; Wu et al., 2023a) and proven effectiveness in audio-text learning. For the HT-SAT audio encoder, all audio clips are randomly cropped or padded to 10 seconds to meet HT-SAT’s fixed-size input requirement.
@@
4.2. Baselines
As we are the first to introduce the task of Text-Audio Incremental Learning and there is no prior related work, we selected the state-of-the-art method in the field of audio incremental learning AI-CL (Muthuchamy Selvaraj et al., 2023) and several effective prompt tuning approaches, VPT (Jia et al., 2022) and I-VL (Ju et al., 2022b), to build a benchmark. We also selected another popular parameter-efficient transfer learning approach LoRA (et al., 2022), which learns low-rank matrices to approximate parameter updates and reduce the number of trainable parameters. These methods have been shown to be effective in fine-tuning models and enabling efficient transfer across various tasks. Meanwhile, we also compare our method with the fine-tuning method.
@@ 4.2. Baselines
  • AI-CL (Muthuchamy Selvaraj et al., 2023). We use the Audio Spectrogram Transformer with Frequency-Time factorized Attention, as proposed in their method, as our audio encoder. In Step 1, we fully train both the audio encoder and the text encoder; in subsequent steps, we follow their approach. Specifically, we utilize the proposed adapter, training only the adapter while freezing the rest of the structure. The model’s input and training setup are the same as in our method.
  • VPT (Jia et al., 2022). We compared the two types of prompts discussed in the paper: (1) VPT (Shallow) prompts and (2) VPT (Deep) prompts. For the shallow prompt approach, we add an audio prompt before the input to the audio encoder. In contrast, for the deep prompt approach, we add a learnable prompt at each layer of the audio encoder. During training, we only train the prompts and projection layers while freezing the rest of the model.
  • I-VL (Ju et al., 2022b). We incorporate a text prompt into our text embeddings, while the audio embeddings remain the same. Similar to VPT, we freeze the audio and text encoder, only training the text prompts and projection layers.
  • LoRA (et al., 2022). To apply LoRA, we applies trainable low-rank updates to $W_{q}$ and $W_{v}$ in Multi-Head Self Attention, and we’ve integrated this into the audio encoder, freezing its other parameters and the text encoder.
    Figure 4. Average retrieval performance across four datasets
  • Finetuning approach. We designed two variations: (1) Finetune (Sequential). Follows a sequential finetuning process on each dataset, aligning with incremental learning. (2) Finetune (Joint). Employs a joint training approach, combining all datasets. In the two finetuning methods, all parameters are trained without additional prompts or knowledge distillation, effectively using the basic contrastive audio-text learning approach from (Elizalde et al., 2023).
@@
(a) Step Results for AudioCaps dataset. (AudioCaps is introduced at Step 1.)
(b) Step Results for Clotho dataset. (Clotho is introduced at Step 2.)
Figure 5. Incremental step evaluation on AudioCaps (Step 1) and Clotho (Step 2).
Table 2. Anti-Forgetting score on AudioCaps and Clotho.
Method AudioCaps Clotho
AI-CL (Muthuchamy Selvaraj et al., 2023) 0.44 0.80
VPT (Shallow) (Jia et al., 2022) 0.65 0.83
VPT (Deep) (Jia et al., 2022) 0.43 0.72
I-VL (Ju et al., 2022a) 0.66 0.84
LoRA (et al., 2022) 0.79 0.94
Finetune (Sequential) 0.61 0.68
PTAT 0.85 0.92
Table 3. Ablation study on the effect of adding audio prompts at different layers of the audio encoder.
@@
4.3. Main Results
We present our main experimental results after sequential training on four datasets, with the training order being AudioCaps(Step 1), Clotho (Step 2), BBC Sound Effects (Step 3), and AudioSet (Current Step). As shown in Tab. 1, our method PTAT significantly outperforms AI-CL, VPT (Shallow), VPT (Deep), and I-VL in the 1-3 steps, specifically AudioCaps (Step 1) and Clotho (Step 2). In the Current Step, we maintain a comparable level with these four methods in both the Audio-to-Text and Text-to-Audio sub-tasks across $R@1$, $R@5$, and $R@10$ metrics. This demonstrates that our method more effectively prevents catastrophic forgetting and exhibits superior model generalization capabilities.
Table 4. Ablation study on different feature and similarity distillation loss items in our method. Bold and underline indicate the best and the second-best results.
@@ 4.3. Main Results
$\mathcal{L}_{FD}$ $\mathcal{L}_{SD}$ R@1 R@5 R@10 R@1 R@5 R@10
11.52+0.00 27.10+0.00 35.15+0.00 16.33+0.00 36.74+0.00 47.53+0.00
12.42+0.90 28.78+1.68 37.02+1.87 18.19+1.86 38.24+1.50 50.50+2.97
13.02+1.50 29.33+2.23 38.15+3.00 18.14+1.81 39.24+2.50 51.11+3.58
13.10+1.58 29.69+2.59 38.11+2.96 19.30+2.97 40.35+3.61 50.61+3.08
Figure 6. Ablation study on the effect of different dataset introduction orders.
@@
To provide a more comprehensive comparison of model performance, we use the average recall across these four datasets as the final evaluation metric for each method. As shown in Fig. 4, in the Text-to-Audio and Audio-to-Text retrieval tasks, our method outperforms other methods on $R@1$, $R@5$, and $R@10$ metrics. Among methods with similar parameter sizes, including VPT (Shallow), VPT (Deep), and I-VL, I-VL achieves the best results. Compared to the I-VL method with a similar parameter scale, our method achieves an average improvement of 7% on average. Due to the domain gaps across datasets, the Finetune (Joint) method, despite using a larger parameter size, performs significantly worse than the Finetune (Sequential) method. Compared to the large-scale Finetune (Sequential) method, our method only requires 2.42% of its parameters, while still achieving 4.46% higher performance.
To more intuitively demonstrate the model’s resistance to forgetting, as shown in Fig. 5, we focus on the model’s memory performance on the first two datasets. Fig. 5(a) illustrates the incremental learning performance on AudioCaps (introduced at Step 1), while Fig. 5(b) shows the test accuracy of Clotho (introduced at Step 2) at each incremental step. It is evident that our PTAT method experiences the smallest accuracy drop in each incremental step and consistently maintains the highest accuracy, demonstrating strong resistance to forgetting and stable accuracy throughout the training process. Additionally, we observed that VPT (Deep) exhibits more severe forgetting compared to VPT (Shallow), suggesting that due to the relatively simple structure of the audio encoder, excessive prompts may lead to more pronounced forgetting.
These trends are further supported by the anti-forgetting scores reported in Tab. 2, which quantify the degree of knowledge retention. Our method consistently outperforms the baselines on both datasets. Notably, the score on AudioCaps is slightly lower than that on Clotho, given that AudioCaps was introduced earlier and thus more susceptible to forgetting due to a longer exposure to subsequent training. These findings collectively highlight the robustness and effectiveness of our approach in mitigating forgetting across different stages of incremental learning.
@@
4.4. Ablation Studies
RQ 1: How does adding a prompt layer affect model performance in TAIL tasks?
@@ 4.4. Ablation Studies
To analyze the different effects of adding prompts at various layers within the encoder, we conducted experiments on different transformer layers of the audio encoder, with the results shown in Tab. 3. Since HTSAT only has 4 layers, we conducted experiments on the first, second, and third layers. We have found that adding prompts at Layer 1 significantly enhances the model’s resistance to forgetting, particularly achieving great performance on the AudioCaps and Clotho datasets. It also requires fewer parameters than adding prompts at Layer 2 and Layer 3. In contrast, adding prompts at Layer 3 helps the model learn new datasets, as evidenced by improved results on the last two datasets, BBC Sound Effects and AudioSet. However, it reduces the model’s ability to retain information from historical datasets.
RQ 2: How do the different feature and similarity distillation loss components impact the performance of our method?
@@
We analyze the effectiveness of our proposed Audio-Text Similarity and Feature Distillation module by removing single or both parts from our PTAT model. As shown in Tab. 6, using Similarity and Feature Distillation (Full ($\mathcal{L}$)) achieves the best performance compared to no distillation or using only individual Similarity/Feature Distillation.This demonstrates the effectiveness of our proposed Similarity and Feature Distillation module.Among them, the combination of removing both Similarity Distillation ($\mathcal{L}_{SD}$) and Feature Distillation ($\mathcal{L}_{FD}$) performs the worst. Compared to adding only Feature Distillation, adding only Similarity Distillation results in a more significant improvement.
RQ 3: Does the order of incremental learning in the dataset affect the experimental results?
To simulate real-world data dynamics, we evaluated our model’s performance under different dataset introduction orders. Specifically, AudioCaps was introduced at Step 1, followed by Clotho at Step 2, BBC Sound Effects at Step 3, and AudioSet at Step 4 (the Current Step). As shown in Fig. 6, the results demonstrate that our method remains effective regardless of the dataset order, highlighting its adaptability to flexible and evolving data updates.
5. Conclusion
@@
We introduces TAIL, a novel text-audio incremental learning task. Its challenges focused on achieving continuous text-audio retrieval tasks across sequentially expanding datasets, which is previously unexplored. To benchmark the TAIL task, we propose a parameter-efficient method named PTAT, which tackles the task by learning cross-modal prompts tuning and incorporating an audio-text similarity and feature distillation module. PTAT consists of two components: Audio-Text Prompt Generation (ATPG) and Audio-Text Similarity and Feature Distillation (ATSFD). ATPG generates a series of interconnected audio and text prompts, leveraging prompt tuning to optimize parameters efficiently. ATSFD mitigates catastrophic forgetting by constraining the output of the new model based on the output of the previous model. We benchmark our method and previous incremental learning or prompt tuning methods on AudioCaps, Clotho, BBC Sound Effects and AudioSet datasets, and our method outperforms previous methods significantly, particularly demonstrating stronger resistance to forgetting on older datasets. Notably, compared to the full-parameter Finetune (Sequential) approach, our model achieves a 4.46% performance improvement while using only 2.42% of the parameters, demonstrating strong parameter efficiency and enhanced resistance to catastrophic forgetting.